# FAIR Enough: How Can We Develop and Assess a FAIR-Compliant Dataset for Large Language Model Training?

### 19 Jan 2024 v1

@@ Abstract
Advancements in Large Language Models (LLMs) highlight the need for ethical practices and data integrity. We introduce a framework that embeds FAIR (Findable, Accessible, Interoperable, Reusable) data principles into LLM training. This approach marks a shift towards practices compliant with FAIR standards. Our framework presents guidelines for integrating FAIR data principles into LLM training. This initiative includes a checklist for researchers and developers. We also demonstrate its practical application through a case study focused on bias identification and mitigation in our FAIR-compliant dataset. This work is a significant contribution to AI ethics and data science, advocating for balanced and ethical training methods in LLMs.
@@ 1 Introduction
The FAIR data principles, which stand for Findable, Accessible, Interoperable, and Reusable Dunning et al. (1970); Boeckhout et al. (2018), were initially established to improve the stewardship of scientific data. These principles have gained increasing recognition in the field of responsible AI development Deshpande and Sharp (2022). In recent times, particularly in the context of generative AI, the importance of these principles has become more evident due to ethical challenges. These challenges include concerns about bias, privacy, and potential misuse of AI-generated content of Ethics (2024); Partescano et al. (2021). For LMs, including advanced LLMs  [^1] In this study, ’LLMs’ encompasses both ’LMs’ and ’LLMs’, with the latter representing advanced iterations of the former, adherence to FAIR data principles is even more crucial. The way data is curated and managed significantly affects the reliability, ethical integrity, and impact of these models.
The exploration of FAIR data principles has been pivotal in data science and management. Seminal works Dunning et al. (1970); Wilkinson et al. (2016); Hasnain and Rebholz-Schuhmann (2018); Jacobsen et al. (2020a); Shmueli et al. (2021) have provided extensive insights into these principles across various fields. However, their application in training and developing LLMs is an emerging and evolving area of research. Recent advances focus on aligning LLMs with ethical standards Chang et al. (2023); Wang et al. (2023), underscoring the importance of embedding FAIR principles in LLM training from the start. Specific challenges in LLMs, such as data biases Ntoutsi et al. (2020) and interpretability Singh et al. (2023), highlight the need for ethical and balanced training methods. This emphasis on data ethics is in line with broader trends in AI ethics and data science Raji et al. (2021); Jobin et al. (2019) stress the necessity of integrating these principles in LLM development.
We suggest a lifecycle model for developing LLMs that integrates FAIR principles at every phase. A key element of our methodology is developing a FAIR-compliant dataset that includes a range of narratives from various sources. Our main contribution extends beyond just data collection; it involves meticulously aligning this dataset with FAIR principles to improve LLM training effectiveness. While adhering to FAIR principles doesn’t guarantee optimal ethical compliance, it is a significant step in that direction. This research paves the way for further exploration in this area.The key contributions of our work are:
@@ 1.1 Contributions
    We present a detailed review of FAIR data principles within AI research, including a comprehensive checklist for researchers and developers.
@@
    We introduce a novel framework that embeds FAIR data principles throughout the LLM training lifecycle, facilitating ethical and effective use in various AI contexts.
@@
    We demonstrate the practical benefits of a FAIR-compliant dataset through a case study. This study focuses on identifying and addressing biases [^2] Bias refers to a wide range, however, our case study particularly address linguistic biases, either explicit or implicit, targetting certain protected groups Fletcher et al. (2021); Garrido-Muñoz et al. (2021) before training LLMs.
@@ 1.2 Literature Selection Criteria
This study examines publications from the last five years, prioritized English-language articles from prominent AI and ML journals and conferences. We explored databases and proceedings from Elsevier, Springer, Nature portfolio, IEEE Transactions, and key conferences like NeurIPS and ICML. Recognizing the fast-paced evolution of LLM research, we included relevant preprints and seminal works for a comprehensive perspective.
Search Query:
("FAIR principles" OR "data management") AND "large language models"
AND "training data" AND "ethical considerations" FROM 2018 TO 2023 IN English
IN Journals and Conferences (e.g., JAIR, JMLR, IEEE Transactions, NeurIPS, ICLR).
This search resulted in 135 papers, out of which we carefully chose about 75 that directly pertained to FAIR data principles in the context of training models. We noted a scarcity of work combining LLMs with FAIR data principles, a gap that our study aims to address.
@@ 1.3 Comparative Analysis of Related Work
Yes (BERT/ , GPT families, LLaMa-2-7b)
Table 1: Summary of Review Papers on FAIR Data Principles
Table 1 highlights a predominant focus on health data in existing research. Notably, one study incorporates a LLM application (GPT). Our work, however, stands out for its application of FAIR data principles to a range of LLMs, including BERT-large, GPT, and LLaMa-2-7b. This approach broadens the scope and generalizability of our research.
@@ 2 FAIR Data Principles: Theoretical Background and Significance
The FAIR data principles significantly influence data management practices across various fields. These include physics Dungkek (2022), health Inau et al. (2023), environmental science Partescano et al. (2021), pharmaceuticals Wise et al. (2019), chemistry Jeliazkova et al. (2023), computer science Axton et al. (2016), research and development Wilcox (2018), and clinical studies Vesteghem et al. (2019). These principles ensure data is well-organized, accessible, and reusable in different research domains and applications. In AI and Natural Language Processing (NLP), the implementation of interdisciplinary strategies Huerta et al. (2023) and structured approaches Bernabé et al. (2023) is crucial to uphold ethical data practices Shmueli et al. (2021) and ensure fairness in AI models Bateni et al. (2022). Our detailed literature review elaborates on each FAIR principle, as depicted in Figure 1, and we also provide a comprehensive compliance checklist in Table 2333For hyperlinks to the technologies and APIs mentioned in the table, refer to Appendix .1.
@@
2.1 Findability
Findability, a crucial aspect of the FAIR data principles, ensures data and resources are easily locatable and accessible Findlay and Seah (2020). It involves creating detailed metadata, using persistent identifiers, and promoting efficient data indexing and searchability. This enhances discoverability by both humans and machines. Significant contributions to data findability within the FAIR framework include Bernabé et al. (2023), which explores strategies for FAIRification of data to enhance findability, and Santos et al. (2023), investigating a model for the FAIR digital object framework focused on digital objects’ findability. The application of FAIR principles in AI, particularly the importance of findability in AI datasets, is discussed in Huerta et al. (2023). Furthermore, Götz (2023) considers the trustworthiness of data repositories, emphasizing the critical role of findability in these environments.
2.2 Accessibility
Accessibility, as defined by the FAIR data principles, is the ease of obtaining and using data once located Boeckhout et al. (2018). It involves long-term data preservation, ethical access protocols, and ensuring data remains retrievable and usable over time. Discussions on trust in data repositories as part of accessibility are covered in Wang and Savard (2023), which also provides a comprehensive understanding of FAIR principles, including accessibility. The application of FAIR principles to research software, with a focus on accessibility, is explored in Lamprecht et al. (2020). How the Fedora platform supports FAIR data principles, particularly accessibility, is discussed in Wilcox (2018). Challenges and opportunities in implementing FAIR principles, including accessibility, in Brazilian data science are addressed in Sales et al. (2020).
2.3 Interoperability
Interoperability, as outlined in the FAIR data principles, refers to the ability of different data systems to work together efficiently da Silva Santos et al. (2016). It requires standardized data formats and protocols, enabling easy data exchange and integration across diverse systems. The Immune Epitope Database (IEDB) and its focus on interoperability in implementing FAIR principles are discussed in Vita et al. (2018). The eXtensible ontology development (XOD) principles and tools, which support interoperability in ontology development, are explored in He et al. (2018). Challenges and strategies for implementing FAIR data principles, including interoperability, in data science are addressed in Sales et al. (2020). A framework and metrics for evaluating the FAIRness of data, with a focus on interoperability, are provided in Wilkinson et al. (2018). The GO FAIR initiative, which aims to accelerate the adoption of FAIR principles including interoperability, is discussed in Schultes et al. (2018).
2.4 Reusability
Reusability, a key component of the FAIR data principles, ensures data is stored and documented for future retrieval and reuse Anguswamy and Frakes (2012). It encompasses rich metadata, legal and ethical considerations, and potential societal impact. A structured approach to planning data FAIRification, with an emphasis on reusability, is discussed in Bernabé et al. (2023). A model for digital objects under FAIR principles, focusing on reusability, is offered in Santos et al. (2023). The application of FAIR principles in AI, particularly regarding the reusability of AI datasets, is explored in Huerta et al. (2023). The importance of designing workflows with reusability as a primary goal is discussed in Wolf et al. (2021). A comprehensive exploration of FAIR principles, including detailed insights on reusability, is presented in Jacobsen et al. (2020a).
@@ 3 Data Management Challenges in Large Language Models
Understanding the Context: Enhanced NLP algorithms are necessary to bridge the current gap in contextual awareness Xi et al. (2023). The capacity for models to capture and interpret nuanced context is essential for advancing the intelligence and applicability of LLMs.
Accuracy and Reliability: The occasional delivery of factually incorrect responses or hallucinations Ji et al. (2023) in LLMs highlights the need for improved fact-checking mechanisms. Ensuring the accuracy and reliability of responses is crucial for the trustworthiness of these models.
Ethical and Fair Use: To combat potential biases in responses, there is a pressing need for advanced bias detection algorithms Deshpande and Sharp (2022); Wang et al. (2023); Raza et al. (2022). Ensuring the ethical application of LLMs requires a steadfast commitment to fairness and equity in model responses.
Interactivity and Personalization: Current LLMs sometimes provide generic and non-personalized responses. To address this, there is a need for adaptive learning algorithms that can learn from user interactions and provide personalized experiences He et al. (2023).
Language and Cultural Sensitivity The limited support for low-resource languages indicates a need for expanded language models and cultural datasets Chang et al. (2023). This will promote inclusivity and cultural sensitivity in LLMs.
Efficiency and Scalability: A significant gap in LLM development is the high computational resource requirement Zhao et al. (2023b). There is a need for optimization strategies that enable efficient processing without compromising the model’s performance.
@@
: The challenges of managing data for LLMs are multifaceted, but the FAIR data principles can offer some targeted solutions. For example, Findability addresses the complexities of vast datasets by enhancing metadata and providing persistent identifiers. Interoperability standardizes data formats, enabling reliable annotation and labeling. It also plays a role in contextual awareness, improving the capacity of NLP algorithms to understand nuanced contexts. Accessibility further governs the delicate balance between sharing data and protecting proprietary information, while reusability confirms that data practices are compliant with legal and regulatory standards, ensuring that data remains usable and trustworthy. We summarize these suggestions to mapping data management challenges in LLMs to FAIR data principles in Appendix .4.
@@ 4.1 Rationale for Integrating FAIR Data Principles in LLM Development
Our research focuses on improving Large Language Models (LLMs) by incorporating FAIR data principles into their training methods. We aim to create training datasets that adhere to FAIR principles specifically designed for LLMs. While FAIR principles have been a standard practice in AI Liang et al. (2021); Navigli et al. (2023) for a while, their integration into LLMs is a novel contribution from our team. To achieve this goal, we’ve developed a comprehensive framework that integrates FAIR principles across the entire lifecycle of LLMs, as illustrated in Figure 3. Next, we detail each phase of the LLM lifecycle, linking it specifically to our case study focused on the development of a FAIR-compliant dataset.
@@
4.2 Case Study: Developing a Bias-Identification Dataset
To address the risk of LLMs perpetuating societal biases, our case study focuses on developing a dataset with a conscientious design that proactively identifies biases within the data prior to the training of these models ( whether they be LLMs or other ML models).
In this case study, “bias" is identified as a tendency within the dataset leading to the unfair representation of specific groups or outcomes Ntoutsi et al. (2020). This bias may take the form of linguistic bias Raza et al. (2024), where the dataset could unfairly favor or disfavor certain protected groups, identified by race, gender, or ethnicity, in terms of languages or dialects. It might also manifest as stereotypes Nadeem et al. (2021), toxicity Barikeri et al. (2021), or misinformation Raza and Ding (2022). Such biases typically originate from the methods used in data collection, processing, or usage. As a result, they can cause these LLMs to replicate or magnify such biases May et al. (2019). Incorporating FAIR data principles into the creation of LLMs is a proactive measure to identify and mitigate these biases before they are amplified in model outcomes. The diversity of bias dimensions and aspects addressed in our dataset is outlined in Figure 4, which shows a broad spectrum of biases including ageism, occupational jargon barriers, political rhetoric, and more.
The primary objective of this case study, aside from bias identification, is to establish a model dataset that adheres to FAIR principles. This dataset can serve as a template or example for creating datasets with various objectives for training LLMs while ensuring compliance with FAIR principles.

### 23 Jan 2024 v2

# FAIR Enough: Develop and Assess a FAIR-Compliant Dataset for Large Language Model Training?

@@ Abstract
The rapid evolution of Large Language Models (LLMs) underscores the critical importance of ethical considerations and data integrity in AI development, emphasizing the role of FAIR (Findable, Accessible, Interoperable, Reusable) data principles. While these principles have long been a cornerstone of ethical data stewardship, their application in LLM training data is less prevalent, an issue our research aims to address. Our study begins with a review of existing literature, highlighting the significance of FAIR principles in data management for model training. Building on this foundation, we introduce a novel framework that incorporates FAIR principles into the LLM training process. A key aspect of this approach is a comprehensive checklist, designed to assist researchers and developers in consistently applying FAIR data principles throughout the model development lifecycle. The practicality and effectiveness of our framework are demonstrated through a case study that involves creating a FAIR-compliant dataset to detect and reduce biases. This case study not only validates the usefulness of our framework but also establishes new benchmarks for more equitable, transparent, and ethical practices in LLM training. We offer this framework to the community as a means to promote technologically advanced, ethically sound, and socially responsible AI models.
@@ 1 Introduction
The FAIR data principles, which stand for Findable, Accessible, Interoperable, and Reusable Dunning et al. (1970); Boeckhout et al. (2018), were initially established to improve the stewardship of scientific data. These principles are highly beneficial for any model development lifecycle Wise et al. (2019) and have become increasingly recognized in responsible AI development Deshpande and Sharp (2022). Recently, the relevance of FAIR principles has been particularly underscored in generative AI due to ethical challenges such as bias, privacy concerns, and the potential misuse of AI-generated content of Ethics (2024); Partescano et al. (2021). This highlights the growing importance of ensuring data used in AI is findable, accessible, interoperable, and reusable, adhering to ethical standards. For LMs, including advanced LLMs 111’LLMs’ here includes both ’LMs’ and ’LLMs’, with ’LLMs’ representing more advanced versions, adherence to FAIR data principles is becoming crucial. The manner in which data is curated and managed significantly impacts the reliability, ethical integrity, and overall effectiveness of these models.
Seminal works in data science and management Dunning et al. (1970); Wilkinson et al. (2016); Hasnain and Rebholz-Schuhmann (2018); Jacobsen et al. (2020); Shmueli et al. (2021) have explored FAIR principles in various domains. Their application in the training and development of LLMs, however, is a developing area of research. While recent advances in LLM studies Chang et al. (2023); Wang et al. (2023) have focused on aligning LLMs with ethical standards, the direct incorporation of FAIR principles in LLM training is less explicit. Challenges specific to LLMs, such as addressing data biases Ntoutsi et al. (2020) and enhancing model explainability and interpretability Singh et al. (2023), further highlight the need for ethically and balanced training approaches that implicitly align with FAIR principles. This emphasis on data ethics is in line with broader trends in AI ethics and data science Raji et al. (2021); Jobin et al. (2019) stress the necessity of integrating these principles in LLM development.
In this work, we shed some light on FAIR data principles and propose a model development lifecycle for LLM training that meticulously incorporates FAIR principles at each stage. At the heart of our approach is the creation of a FAIR-compliant dataset, designed to include a wide array of narratives from diverse sources. The central contribution of our work, however, extends beyond simple data collection and preparation for LLM training. We have placed a strong emphasis on rigorously aligning the dataset with FAIR principles, significantly boosting the training efficiency of LLMs. We acknowledge that adherence to FAIR principles (even at the strict-most level) may not equate to absolute ethical compliance, however, it represents a crucial step in that direction. Our research establishes a foundational framework for further advanced studies in this field. The primary contributions of our work are:
@@ 1.1 Contributions
    An exploration of FAIR data principles within the realm of AI research, including the provision of a comprehensive checklist for researchers and developers.
@@
    The introduction of an innovative framework that integrates FAIR data principles throughout the LLM training lifecycle, ensuring ethical and effective application in various AI contexts.
@@
    The demonstration of the practical benefits of a FAIR-compliant dataset through a case study. This study specifically focuses on identifying and mitigating biases  [^2] The topic of bias is broad-ranging, our case study specifically focuses on addressing linguistic biases targeting protected groups. Fletcher et al. (2021); Garrido-Muñoz et al. (2021) prior to training LLMs. This approach utilizes FAIR principles in building the dataset and in training various LLMs.
The remainder of the paper is structured as follows: Section 2 presents “FAIR Data Principles: Theoretical Background and Significance". In Section 3, we delve into “Data Management Challenges in Large Language Models". Section 4 is dedicated to “Framework for Integrating FAIR Data Principles in the Development of Large Language Models". The paper then moves to Section 5 for a comprehensive “Discussion", and concludes with the Section 6 for “Conclusion".
@@ 1.2 Literature Selection Criteria
This study examines publications from the last five years, in English-language articles from prominent AI and ML journals and conferences. We explored several databases and proceedings from high-quality journals from Elsevier, Springer, Nature portfolio, IEEE Transactions, and key conferences like NeurIPS, ICML, ACL and such. Recognizing the fast-paced evolution of LLM research, we included relevant preprints and seminal works for a comprehensive perspective. This search resulted in 135 papers, out of which we carefully chose about 75 that directly pertained to FAIR data principles in the context of training models. We noted a scarcity of work combining LLMs with FAIR data principles, a gap that our study aims to address.
The search query is :
("Findability" OR "data discovery" OR "metadata standards" OR "persistent identifiers") AND ("Accessibility" OR "data access" OR "data sharing policies" OR "authentication and authorization mechanisms") AND ("Interoperability" OR "data integration" OR "standardized data formats" OR "cross-domain data exchange") AND ("Reusability" OR "data documentation" OR "data quality assurance" OR "long-term data preservation") AND ("large language models"
OR "LLMs" OR "AI models" OR "machine learning models") AND ("training data management" OR "ethical data sourcing" OR "bias in AI datasets" OR "responsible data use in AI") AND ("ethical considerations in AI" OR "AI ethics" OR "responsible AI" OR "ethical AI development") PUBLISHED FROM 2018 TO 2023 IN English IN Journals and Conferences (e.g., JAIR, JMLR, IEEE Transactions, NeurIPS, ICLR, ACL, Springer, Nature portfolio, Elsevier).
@@ 1.3 Comparative Analysis of Related Work
Table 1: Summary of Review Papers on FAIR Data Principles
@@
Yes (BERT , GPT families, LLaMa-2-7b)
Table 1 highlights a focus on health data in existing research on FAIR principles. Notably, one study incorporates a LLM application (GPT-based model). Our work, however, stands out for its application of FAIR data principles to a range of LLMs, including BERT-large, GPT, and LLaMa-2-7b. This approach broadens the scope and generalizability of our research.
@@ 2 FAIR Data Principles: Theoretical Background and Significance
The FAIR data principles significantly influence data management practices across various fields. These include physics Dungkek (2022), health Inau et al. (2023), environmental science Partescano et al. (2021), pharmaceuticals Wise et al. (2019), chemistry Jeliazkova et al. (2023), computer science Axton et al. (2016), research and development Wilcox (2018), and clinical studies Vesteghem et al. (2019). These principles ensure data is well-organized, accessible, and reusable in different research domains and applications. In AI and Natural Language Processing (NLP), the implementation of interdisciplinary strategies Huerta et al. (2023) and structured approaches Bernabé et al. (2023) is crucial to uphold ethical data practices Shmueli et al. (2021) and ensure fairness in AI models Bateni et al. (2022). Given the constraints of space, we present a concise overview of each FAIR principle, as depicted in Figure 1. Additionally, a compliance checklist is provided in Table 2333For hyperlinks to the technologies and APIs referenced in the table, refer to Appendix .1.
@@
2.1 Findability
Findability, a crucial aspect of the FAIR data principles, ensures data and resources are easily locatable and accessible Findlay and Seah (2020). It involves creating detailed metadata, using persistent identifiers, and promoting efficient data indexing and searchability. This enhances discoverability by both humans and machines. Some contributions to data findability within the FAIR framework include Bernabé et al. (2023), which explores strategies for FAIRification of data to enhance findability, and Santos et al. (2023), investigating a model for the FAIR digital object framework focused on digital objects’ findability. The application of FAIR principles in AI, particularly the importance of findability in AI datasets, is discussed in Huerta et al. (2023). Furthermore, Götz (2023) considers the trustworthiness of data repositories, emphasizing the critical role of findability in these environments.
2.2 Accessibility
Accessibility, as defined by the FAIR data principles, is the ease of obtaining and using data once located Boeckhout et al. (2018). It involves long-term data preservation, ethical access protocols, and ensuring data remains retrievable and usable over time. Discussions on trust in data repositories as part of accessibility are covered in Wang and Savard (2023), which also provides a comprehensive understanding of FAIR principles, including accessibility. The application of FAIR principles to research software, with a focus on accessibility, is explored in Lamprecht et al. (2020). How the Fedora platform supports FAIR data principles, particularly accessibility, is discussed in Wilcox (2018). Challenges and opportunities in implementing FAIR principles, including accessibility, in Brazilian data science are addressed in Sales et al. (2020). These works emphasize the significance of incorporating accessibility considerations into research software, platform support, and implementation challenges.
2.3 Interoperability
Interoperability, as outlined in the FAIR data principles, refers to the ability of different data systems to work together efficiently da Silva Santos et al. (2016). It requires standardized data formats and protocols, enabling easy data exchange and integration across diverse systems. The Immune Epitope Database (IEDB) and its focus on interoperability in implementing FAIR principles are discussed in Vita et al. (2018). The eXtensible ontology development (XOD) principles and tools, which support interoperability in ontology development, are explored in He et al. (2018). Challenges and strategies for implementing FAIR data principles, including interoperability, in data science are addressed in Sales et al. (2020). A framework and metrics for evaluating the FAIRness of data, with a focus on interoperability, are provided in Wilkinson et al. (2018). The GO FAIR initiative, which aims to accelerate the adoption of FAIR principles including interoperability, is discussed in Schultes et al. (2018). All these works underscore the role of interoperability in advancing responsible and efficient data management in various scientific and technological fields.
2.4 Reusability
Reusability, a key component of the FAIR data principles, ensures data is stored and documented for future retrieval and reuse Anguswamy and Frakes (2012). It encompasses rich metadata, legal and ethical considerations, and potential societal impact. A structured approach to planning data FAIRification, with an emphasis on reusability, is discussed in Bernabé et al. (2023). A model for digital objects under FAIR principles, focusing on reusability, is offered in Santos et al. (2023). The application of FAIR principles in AI, particularly regarding the reusability of AI datasets, is explored in Huerta et al. (2023). The importance of designing workflows with reusability as a primary goal is discussed in Wolf et al. (2021). All these works emphasize that data should not only be archived for future use but also be ethically and efficiently usable, contributing long-term value to the scientific and technological data ecosystem.
An exploration of FAIR principles, including detailed insights on reusability, is presented in Jacobsen et al. (2020).
@@ 3 Data Management Challenges in Large Language Models
Understanding the Context: Enhanced NLP algorithms are vital for improving contextual awareness in LLMs Xi et al. (2023). The ability of these models to discern and interpret subtle and complex contexts is key in elevating their intelligence and practical applicability. There is a growing need for algorithms capable of understanding linguistic subtleties, cultural nuances, and the varied implications of context in human language.
Accuracy and Reliability: The issue of LLMs occasionally generating factually incorrect responses or experiencing ’hallucinations’ Ji et al. (2023) underscores the importance of robust fact-checking mechanisms. Enhancing the trustworthiness and truthfulness of these models requires advanced verification algorithms and data validation methods to ensure accuracy and reliability.
Ethical and Fair Use: Addressing inherent biases in LLM responses is critical Deshpande and Sharp (2022); Wang et al. (2023); Raza et al. (2022). The development of advanced bias detection algorithms is crucial, involving systems that identify and mitigate biases in training data and model responses, promoting fairness and equity Raza et al. (2023). This requires a concerted effort to understand and reflect diverse perspectives, ensuring that LLMs are used ethically, promoting fairness and equity in their outputs.
Interactivity and Personalization: The tendency of current LLMs to provide generic responses points to the need for more sophisticated adaptive learning algorithms He et al. (2023). These algorithms should be capable of learning from individual user interactions, preferences, and feedback, thus offering increasingly personalized and relevant responses. The goal is to develop models that can understand and adapt to individual user needs, enhancing the overall user experience through personalization.
Language and Cultural Sensitivity: The limitations in supporting low-resource languages call for the expansion of language models and cultural datasets Chang et al. (2023). Enhancing the representation of diverse languages and cultures in LLMs is essential for promoting global inclusivity and cultural sensitivity . This involves not only expanding the datasets with more languages but also ensuring that these models understand and respect cultural nuances and contexts.
Efficiency and Scalability: A major challenge in LLM development is the high computational resource requirement Zhao et al. (2023b). To address this, it is necessary to develop and implement optimization strategies that maintain or improve the performance of LLMs while significantly reducing their computational demands. This could involve innovations in model architecture, data processing techniques, and hardware optimizations, enabling more efficient and scalable deployment of LLMs.
@@
: These above-summarized challenges of managing data for LLMs are multifaceted, but the FAIR data principles can offer some targeted solutions. For example, Findability not only simplifies the task of locating relevant data in extensive datasets but also aids in streamlining the research process. Enhanced metadata and persistent identifiers improve the efficiency of data retrieval and analysis, crucial for developing sophisticated LLMs. Interoperability ensures that different data formats and systems can work together seamlessly, which is critical for integrating varied data sources into LLM training.It also supports the development of more contextually aware NLP algorithms by enabling the synthesis of diverse data types and sources, thus contributing to more nuanced and accurate model outputs.
Accessibility involves not just the availability of data but also the mechanisms of access control, ensuring that data is shared responsibly. It balances the need for open data to promote innovation and collaboration against the need to protect sensitive information and intellectual property, a key concern in AI development. Reusability extends the value of data beyond its initial purpose. By ensuring that data practices comply with legal and regulatory standards, it guarantees that data remains relevant, reliable, and ethical for future use. This is particularly important in LLM development, where data may be reused for multiple models or purposes over time. We provide a mapping of these data management challenges in LLMs to FAIR data principles in Appendix .4.
@@ 4.1 Rationale for Integrating FAIR Data Principles in LLM Development
Our research is dedicated to enhancing the training of LLMs by embedding FAIR data principles directly into their training methodologies. The central objective is to develop training datasets for LLMs that are designed to comply with FAIR principles. To realize this, we have devised a framework that interweaves FAIR principles throughout the entire model development lifecycle of LLMs, as depicted in Figure 3. This approach is not limited to just data preparation but extends to every stage of the LLM development process. In the following sections, we will elaborate on each phase of this proposed LLM lifecycle, connecting each step specifically to our case study which concentrates on the creation of a FAIR-compliant dataset.
@@
4.2 Case Study: Developing a FAIR-Compliant Dataset
To address the risk of LLMs perpetuating societal biases, our case study focuses on developing a dataset with a conscientious design that proactively identifies biases within the data prior to the training of these models. The dataset and model card are available here  [^4] Anonymous due to double blind review.
In this case study, “bias" is identified as a tendency within the dataset leading to the unfair representation of specific groups or outcomes Ntoutsi et al. (2020). This bias may take the form of linguistic bias Raza et al. (2024), where the dataset could unfairly favor or disfavor certain protected groups, identified by race, gender, or ethnicity, in terms of languages or dialects. It might also manifest as stereotypes Nadeem et al. (2021), toxicity Barikeri et al. (2021), or misinformation Raza and Ding (2022). Such biases typically originate from the methods used in data collection, processing, or usage. As a result, they can cause these LLMs to replicate or magnify such biases May et al. (2019). We consider a diversity of bias dimensions and aspects in our case study, as outlined in Figure 4, which shows a broad spectrum of biases including ageism, occupational jargon barriers, political rhetoric, and more. This dataset can serve as a template or example for creating datasets with various objectives for training LLMs while ensuring compliance with FAIR principles.